OpenStudy (anonymous):

Translate into an equation and solve. (Use n to represent the smaller number.) The sum of two numbers is three. The difference between nine and twice the smaller number is seven less than four times the larger. Find the numbers Smaller ________________ Larger_________________

OpenStudy (anonymous):

\[\begin{cases}n+m=3\\9-2n+7=4m\end{cases}\]

OpenStudy (anonymous):

so what is the bigger number and smaller number?

OpenStudy (anonymous):

solve this equations and you will get answer, n is smaller number, m is bigger